OpenStudy (highschoolmom2010):

Write the standard equation of each circle. center (0, 3); r =7

OpenStudy (highschoolmom2010):

\[(x-h)^2+(y-k)^2=r^2\] that is the equation for a circle

OpenStudy (isaiah.feynman):

you know it. Now substitute known values.

OpenStudy (highschoolmom2010):

im confused on where to place the (0,3) in the equation

OpenStudy (isaiah.feynman):

(h,k) is the centre of the circle.

OpenStudy (highschoolmom2010):

oh ok

OpenStudy (highschoolmom2010):

\[(x-0)^2+(y-3)^2=7^2\] like that?

OpenStudy (isaiah.feynman):

Yep.

OpenStudy (highschoolmom2010):

x^2+y^2+9=49 ok so far?

OpenStudy (isaiah.feynman):

Can be simplified to